### v1

# Large-scale moral machine experiment on large language models

Abstract
The rapid advancement of Large Language Models (LLMs) and their potential integration into autonomous driving systems necessitates understanding their moral decision-making capabilities. While our previous study examined four prominent LLMs using the Moral Machine experimental framework, the dynamic landscape of LLM development demands a more comprehensive analysis. Here, we evaluate moral judgments across 51 different LLMs, including multiple versions of proprietary models (GPT, Claude, Gemini) and open-source alternatives (Llama, Gemma), to assess their alignment with human moral preferences in autonomous driving scenarios. Using a conjoint analysis framework, we evaluated how closely LLM responses aligned with human preferences in ethical dilemmas and examined the effects of model size, updates, and architecture. Results showed that proprietary models and open-source models exceeding 10 billion parameters demonstrated relatively close alignment with human judgments, with a significant negative correlation between model size and distance from human judgments in open-source models. However, model updates did not consistently improve alignment with human preferences, and many LLMs showed excessive emphasis on specific ethical principles. These findings suggest that while increasing model size may naturally lead to more human-like moral judgments, practical implementation in autonomous driving systems requires careful consideration of the trade-off between judgment quality and computational efficiency. Our comprehensive analysis provides crucial insights for the ethical design of autonomous systems and highlights the importance of considering cultural contexts in AI moral decision-making.
Introduction
The rapid advancement of Large Language Models (LLMs) has inaugurated a new era in artificial intelligence (AI), demonstrating unprecedented capabilities across diverse domains [1, 2, 3]. These models, exemplified by ChatGPT [4, 5] (developed by OpenAI) and its counterparts, exhibit remarkable proficiency in natural language processing, content generation, complex problem-solving, and decision-making. As LLMs evolve, their potential integration into critical systems has become a subject of intense scrutiny and research [6, 7, 8].
Autonomous driving represents a field where LLM application holds particular promise while raising significant ethical questions [9, 10, 11, 12]. The automotive industry is actively exploring LLM integration to enhance autonomous vehicles’ capabilities, potentially improving environmental comprehension, complex scenario interpretation, nuanced decision-making, and communication with passengers and other road users [13, 14, 15, 16, 17, 18, 19]. This integration spans various aspects of autonomous driving, including perception, navigation, planning, and control.
The Moral Machine experiment [10] emphasized the importance of ethical considerations in AI decision-making for autonomous systems. This study presented participants with scenarios involving unavoidable accidents to assess public opinion on how autonomous vehicles should act in morally challenging situations. The experiment provided a framework for exploring human moral preferences, identifying trends such as prioritizing human lives over animals and favoring the preservation of a greater number of lives. While acknowledging its limitations [20, 21, 22, 23, 24, 25], this work has stimulated extensive research into the ethical implications of AI in autonomous systems [26, 27, 28, 29].
While some studies have explored AI responses to standard ethical dilemmas [30, 31], such as the classic trolley problem [32], the intricate scenarios presented in the Moral Machine experiment offer a more comprehensive exploration of AI moral preferences. Building on this approach, our previous study [33] applied the Moral Machine framework to a limited set of LLMs, specifically focusing on ChatGPT (including both GPT-3.5 and GPT-4 versions), Llama 2, and PaLM 2. This investigation examined the moral judgments made by these AI systems in complex autonomous driving scenarios. The results indicated that LLMs were capable of producing ethical judgments that often aligned with human preferences in many aspects. However, we also observed some notable discrepancies between LLM outputs and human choices, particularly in scenarios involving multiple ethical considerations
Despite these initial findings, the rapid evolution of LLM technology necessitates a more comprehensive analysis. This dynamic landscape is evidenced by the frequent release and update of both proprietary models like ChatGPT, Claude [34] (Anthropic), and Gemini [35] (Google), as well as open-source alternatives such as Llama 3 [36] (Meta) and Gemma [37] (Google). The emergence of these diverse models underscores the importance of examining the consistency and evolution of moral judgments across different LLM versions, architectures, and development approaches.
This study aims to conduct a more extensive examination of moral judgment trends in the latest LLM models and investigate the effects of model updates on these judgments. By comparing a wider array of LLMs, we seek to provide insights into the approaches different LLMs take to moral dilemmas, how these approaches may change over time, and the potential implications for AI system deployment in critical decision-making roles, particularly in autonomous driving. This study will contribute to a more nuanced understanding of AI outputs in high-stakes scenarios, informing the development and deployment of AI systems in complex, real-world contexts.
Materials and methods
Moral Machine scenario generation
Following the methodology established in our previous study [33], we generated a comprehensive set of Moral Machine scenarios to evaluate LLM ethical decision-making. These scenarios presented dilemmas involving an autonomous vehicle facing an unavoidable accident, requiring a choice between two outcomes (“Case 1” and “Case 2”) with varying ethical implications. The scenarios were designed through a process of constrained randomization, exploring six primary dimensions: species (humans vs. pets), social value (higher vs. lower perceived social status), gender (female vs. male), age (younger vs. older), fitness (physically favored vs. less fit individuals), and utilitarianism (smaller vs. larger group). Additionally, three secondary dimensions were incorporated: interventionism (swerving vs. continuing straight), relationship to the autonomous vehicle (passengers vs. pedestrians), and legal considerations (e.g., adherence to traffic signals). The scenarios were carefully constructed to balance complexity and clarity, ensuring that they effectively probed the ethical reasoning capabilities of the LLMs under study.
Large language models
This study expanded upon our previous study [33] by incorporating a comprehensive set of large language models (LLMs), including the latest versions available. These LLMs were utilized to generate responses to moral machine scenarios.
We initially incorporated the results from [33], which included responses from three proprietary models: GPT-3.5 (gpt-3.5-turbo-0613), GPT-4 (gpt-4-0613), and PaLM 2, as well as one open source model: Llama 2 (7B). Additional models were selected based on their prominence in the field, technological advancements, and public recognition. This selection aimed to provide a representative sample of current LLM technology while exploring variations in ethical reasoning across different model types.
Among the proprietary models, we evaluated multiple versions of OpenAI models, including the snapshot versions of GPT-3.5 [4] and GPT-4 [5] from March 2023 to April 2024, as well as the newer GPT-4o models [38] (May and August 2024 versions), GPT-4o-mini, o1 and o1-mini [39]. Google DeepMind’s Gemini models were also assessed, including Gemini 1.0 Pro [35] and multiple snapshot versions of Gemini 1.5 Pro and Gemini 1.5 Flash [40]. From Anthropic, we evaluated the latest versions of their Claude models, including Claude 3 Haiku, Sonnet, Opus [34], and Claude 3.5 Sonnet [41].
In the open-source domain, we included Meta’s Llama family [36], comprising Llama 3 and Llama 3.1 in 8B and 70B versions, as well as Llama 3.2 in 1B and 3B versions. A comprehensive evaluation of Google’s Gemma model family [37, 42] was conducted, encompassing multiple versions and scales from 2B-it to 27B-it, including the derivative DataGemma RIG 27B-it model [43]. Additional open-source models evaluated in this study included Large Model Systems Organization’s Vicuna v1.5 (7B and 13B) [44], Mistral (7B Instruct v0.2) [45], Mistral-NeMo [46], CohereForAI’s Command R+ [47], and Microsoft’s Phi-3.5 [48] (both MoE-instruct and mini-instruct versions).
For a comprehensive list of all models and their specific versions used in this study, please refer to our GitHub repository (https://github.com/kztakemoto/mmllm).
All models were used with their default parameter settings. The input prompts were consistent with those used in [33], requiring each model to select either “Case 1” or “Case 2” for each scenario.
While most models were evaluated using 50,000 scenarios, following our previous study [33], constraints related to application programming interface (API) usage costs and computational time necessitated the use of fewer scenarios for certain models. The majority of the models were assessed using the full set of 50,000 scenarios. However, a substantial group of models, including the GPT-4 and GPT-4o series, Claude 3 Opus and Claude 3.5 family, all Gemini family, Llama 3 and subsequent versions, OpenAI o1 mini, Mistral-NeMo, Gemma-2 family, DataGemma, Command R+, and Phi-3.5 family, were evaluated using a reduced set of 10,000 scenarios. The OpenAI o1 model, facing even more significant API cost limitations, was evaluated using a further reduced set of 5,000 scenarios.
Data analysis
For analyzing LLM responses to the Moral Machine scenarios, we employed the conjoint analysis framework as described in [49] and implemented in the original study [10] on the Moral Machine experiment and our previous study [33]. This approach allows for a robust, non-parametric identification of causal effects without specific modeling assumptions.
The analysis process involved pre-processing LLM responses and dummy variable coding for scenario attributes, which correspond to the primary and secondary dimensions described in the scenario generation. We then calculated the Average Marginal Component Effect (AMCE) for each attribute. The AMCE quantifies an attribute’s influence on the LLM’s ethical decisions, enabling systematic comparison of moral preferences across different LLM models and versions.
To compare the moral preference patterns between LLMs and humans, we performed comparative analyses using the calculated AMCE values across all nine attributes. The human preference data were obtained from the original Moral Machine experiment [10]. We quantified the alignment between LLM and human preferences by computing the Euclidean distances between their respective AMCE values. Furthermore, to better understand the relationship patterns among different models and human preferences, we employed principal component analysis (PCA) followed by cluster analysis on the AMCE values.
All statistical analyses, including AMCE calculations, distance measurements, PCA, and subsequent analyses were performed using R statistical software (version 4.4.1).
Results
Moral judgments of large language models
Figure 1 presents radar charts characterizing the moral judgment tendencies of various LLMs across nine preferences quantified by AMCE values (see S1 Table. for exact numerical data). These charts compare the preferences across multiple versions and model sizes of GPT-3.5, GPT-4, GPT-4o/o1, Claude, Gemini, Llama, Gemma, and other LLMs against human preferences (see S1 Fig. for detailed differences between each model and human values).
Analysis revealed that many LLMs demonstrated large positive values (¿0.5) for No. Characters and Species dimensions, indicating tendencies similar to human judgments in prioritizing the preservation of more lives and favoring humans over animals. However, some models exhibited either inverse priorities compared to humans or disproportionately strong preferences in the same direction (differences in AMCE values approximately ¿0.3 compared to human values). We elaborate on these findings below, including the effects of model updates.
{adjustwidth}
-2.25in0in
Fig 1: Radar plots of moral preferences across different LLM families. AMCE values indicate preferences: Species ($+$: humans, --: pets), Social Value ($+$: high status, --: low status), Relation to AV ($+$: pedestrians, -: passengers), Number ($+$: more, --: fewer), Law ($+$: lawful, --: unlawful), Intervention ($+$: inaction, --: action), Gender ($+$: female, --: male), Fitness ($+$: fit, --: unfit/obese), Age ($+$: young, --: elderly). Gray-filled areas represent human preferences. Each subplot represents a different model family: (A) GPT-3.5, (B) GPT-4, (C) GPT-4o/o1, (D) Claude, (E) Gemini, (F) Llama, (G) Gemma, and (H) Other models.
GPT-3.5 family
Analysis of the GPT-3.5 family (Fig 1A) revealed interesting changes associated with model updates. While models from June 2023 to January 2024 exhibited similar preference patterns, the March 2023 version demonstrated distinctive judgment tendencies. For instance, regarding Law, the March version showed positive values favoring law-abiding individuals, aligning with human preferences but displaying stronger preferences than humans. However, subsequent versions shifted toward more neutral judgments with values approaching zero.
Similar changes were observed in Gender-related judgments. The March 2023 version showed strong positive values favoring females, but subsequent updates gradually reduced these values, converging toward human judgment levels. Regarding age-related decisions, while humans demonstrated a preference for younger individuals, all GPT-3.5 versions maintained neutral judgments with values near zero.
For Fitness-related judgments, the March and June 2023 versions showed negative values contrary to human preferences, prioritizing overweight individuals over physically fit ones. However, later versions gradually aligned more closely with human judgments. Regarding Relation to AV, all versions consistently showed positive values similar to humans, but with notably higher magnitudes, indicating a stronger preference for pedestrians over passengers. For Species-related judgments, while the March and June 2023 versions showed values comparable to humans, the November 2023 and January 2024 versions exhibited decreased values, maintaining but weakening the tendency to prioritize humans over pets.
GPT-4 family
Analysis of the GPT-4 family (Fig 1B) revealed high consistency across models from March 2023 to January 2024, demonstrating that model updates had limited impact on moral judgments. Across all nine preference categories analyzed, the models exhibited directional alignment with human judgments, though showing stronger preferences in specific categories.
Most notably, in the Species and No. Characters categories, AMCE values approached 1.0, indicating nearly deterministic preferences for saving humans over animals and larger groups over smaller ones. Similarly, in the Gender, Law, and Relation to AV categories, while aligned directionally with human judgments, the models displayed larger AMCE values. Specifically, they showed more pronounced tendencies than humans in prioritizing females over males, law-abiding individuals over lawbreakers, and pedestrians over passengers.
Conversely, in the Fitness category, the models demonstrated trends diverging from human judgments. While humans showed a slight preference for physically fit individuals, the GPT-4 family exhibited a minor inverse preference, marginally favoring overweight individuals. However, this divergence was relatively minor compared to the pronounced preferences observed in other categories.
GPT-4o/o1 family
Analysis of the GPT-4o family (Fig 1C) similarly demonstrated high consistency across models, with model updates showing limited impact on moral judgments. While this family exhibited AMCE values comparable to human judgments in the Gender and No. Characters categories, notable differences from human preferences were observed across many categories.
Particularly in the Relation to AV category, all models except GPT-4o-mini showed substantially higher values compared to even the GPT-4 family, demonstrating an extreme preference for pedestrians over passengers. The strong preference for law-abiding individuals persisted, and all GPT-4o models showed significantly stronger tendencies than humans in prioritizing humans over pets. Conversely, these models demonstrated relatively weak preferences for younger individuals compared to human age-related judgments, while maintaining a slight, albeit weak, preference for overweight individuals in Fitness-related decisions.
The o1 model generally exhibited judgment patterns similar to the GPT-4o family, though the o1-mini model showed distinctive variations. Notably, o1-mini demonstrated near-neutral preferences aligned with human judgments in the Relation to AV category. However, it recorded AMCE values approaching 1.0 in the No. Characters category, indicating a strong preference for saving larger groups of lives.
Claude family
Analysis of the Claude family (Fig 1D) revealed a distinct bifurcation in judgment tendencies across models. Specifically, two characteristic groups emerged: one comprising Claude 3 Sonnet, Opus, and 3.5 Haiku, and another consisting of Claude 3 Haiku and Claude 3.5 Sonnet.
Claude 3 Sonnet, Opus, and 3.5 Haiku demonstrated distinctive trends that diverged from human judgments across multiple categories. These models showed negative AMCE values in Relation to AV, Fitness, and Social Status categories, indicating inverse preferences compared to humans: prioritizing passengers over pedestrians, overweight over physically fit individuals, and lower over higher social status. While these models aligned directionally with humans in Gender (favoring females over males) and Species (prioritizing humans over pets) categories, the magnitude of their preferences substantially exceeded human values.
Conversely, Claude 3 Haiku and Claude 3.5 Sonnet exhibited different patterns, including neutral age-related judgments that diverged from human preferences. These models also demonstrated excessive preferences, particularly in Relation to AV and Law categories, showing substantially higher AMCE values than humans in prioritizing pedestrians over passengers and law-abiding over law-breaking individuals. Claude 3.5 Sonnet notably demonstrated pronounced tendencies in Species and No. Characters categories, strongly favoring humans over pets and larger over smaller groups.
An interesting finding regarding model evolution revealed that the Sonnet model’s update from 3 to 3.5 generally resulted in closer alignment with human judgments. However, no significant changes in judgment patterns were observed between the June and October versions of Claude 3.5 Sonnet. Conversely, the Haiku model’s update from 3 to 3.5 resulted in greater divergence from human judgments.
Gemini family
Analysis of the Gemini family (Fig 1E) revealed high inter-model consistency across most categories, while demonstrating notable changes in judgment tendencies with updates, particularly in the Relation to AV category. Specifically, while the Gemini 1.0 Pro model showed negative AMCE values indicating preference for passengers over pedestrians, the Gemini 1.5 Pro family shifted to positive values aligned with human judgments, prioritizing pedestrians over passengers.
Similar update-driven changes were observed in the Gemini 1.5 Flash family. While the 1.5 Flash Preview and 1.5 Flash 001 versions prioritized passengers, the update to 1.5 Flash 002 reversed this preference to favor pedestrians. However, the degree of pedestrian preference in the latest versions of both Gemini 1.5 Pro and 1.5 Flash substantially exceeded human values.
Regarding gender-related judgments, only the latest Gemini 1.5 Pro 002 showed values comparable to human judgments, while other versions demonstrated pronounced preferences for females over males. In the Species category, all versions showed stronger preferences than humans for prioritizing humans over pets. Conversely, all models consistently demonstrated inverse tendencies to humans in Fitness-related decisions, showing negative AMCE values favoring overweight over physically fit individuals.
Figure 1E includes PaLM 2, Google’s previous proprietary model, for comparison. PaLM 2 exhibited a distinctive tendency contrary to human judgments, prioritizing smaller over larger groups. In contrast, the Gemini family, Google’s current flagship models, aligned with human judgments across all versions in prioritizing the preservation of more lives.
Llama Family
Analysis of the Llama family (Fig 1F) demonstrated that all models except Llama 3.1 8B showed positive AMCE values similar to humans in the No. Characters category, indicating consistent preferences for saving larger groups of lives.
However, other categories exhibited varying judgment patterns across versions and model sizes. For gender-related decisions, while Llama 2 (7B), 3.2 1B, and 3.2 3B models showed preferences for males contrary to human judgments, other models aligned with humans in favoring females. In the Species category, Llama 3 70B, 3.1 8B, and 3.1 70B models demonstrated clear preferences for humans over pets aligned with human judgments, while other models showed either neutral judgments or slight preferences for pets. Similarly, regarding law compliance, Llama 3 8B, 3 70B, and 3.1 70B models demonstrated preferences for law-abiding individuals, while other models exhibited neutral preferences.
Notably, larger models (70B) showed relatively small variations in judgments between versions and demonstrated closer alignment with human preferences. This observation suggests that model size may influence the consistency of judgment patterns.
Gemma family
Analysis of the Gemma family (Fig 1G) revealed characteristic consistency across several categories. All models demonstrated nearly neutral judgments regarding law compliance and showed consistent preferences for passengers over pedestrians. Notably, these consistent trends, particularly in the pedestrian-passenger prioritization, diverged from human judgments.
However, numerous other categories exhibited varying judgment patterns across models. In the No. Characters category, while Gemma 2 2B and 9B models showed preferences for smaller groups contrary to human judgments, other models aligned with humans in favoring the preservation of more lives. In the Species category, the Gemma 2 series (2B, 9B, 27B) demonstrated preferences for humans over pets aligned with human judgments, while other models showed either neutral judgments or preferences for pets. Regarding Social Status, only Gemma 7B and 1.1 7B aligned with human preferences favoring higher social status, while other models exhibited either neutral judgments or preferences for lower social status.
These variations in judgment consistency and alignment with human preferences suggest potential correlations with model versions and parameter sizes.
Other large language models
Analysis of other LLMs (Fig 1H) generally showed trends similar to human judgments, though with distinctive differences in preference strengths and specific categories. Notably, all models except Vicuna demonstrated higher AMCE values than humans in the Gender category, showing particularly strong preferences for females over males. Additionally, while Mistral and Command R+ aligned with human preferences in favoring larger groups, their preference magnitudes were notably stronger than human values.
Inverse judgment tendencies compared to humans were characteristically observed in the Relation to AV and Fitness categories. Specifically, all models except Vicuna 13B showed negative AMCE values in these categories, demonstrating preferences contrary to human judgments by prioritizing passengers over pedestrians and overweight over physically fit individuals.
Vicuna (7B and 13B) and Phi 3.5 mini exhibited distinctive judgment patterns that generally diverged from human preferences. These models showed neutral judgments regarding law compliance and demonstrated slight preferences for smaller groups, contrary to human tendencies. Particularly, while Vicuna 7B showed neutral judgments in Gender, Age, and Species categories, Vicuna 13B demonstrated characteristic preferences for males over females, pets over humans, and older over younger individuals.
Significantly, different judgment tendencies were observed within the same model families, such as between Vicuna 7B and Vicuna 13B, or between Phi 3.5 mini and Phi 3.5 MoE. These results suggest potential influences of model size on ethical judgments.
Quantitative comparison with human moral judgments
To quantitatively evaluate the similarity between LLM and human moral judgments, we compared the Euclidean distances of AMCE values across nine categories for each model family (Fig 2; see S2 Table. for exact numerical data). This analysis revealed differences between model families and changes in judgment distances from humans through model evolution.
{adjustwidth}
-2.25in0in
Fig 2: Distances between LLMs and human moral judgments across model families. Violin plots show the distribution of distances from human judgments for each model family, with individual models represented as points. Different colors indicate different model families. Model names are labeled.
The GPT-4 family demonstrated the closest alignment with human judgments, with the January 2024 version showing the minimum distance of 0.6. The GPT-3.5 family also exhibited relatively close alignment, with distances ranging from 0.8 to 0.9. However, the newer GPT-4o/o1 family, except for o1-mini, showed a slight increase in distance from human judgments.
In the Claude family, while Claude 3 Haiku and Claude 3.5 Sonnet demonstrated relatively close alignment with human judgments (distances 0.9–1.0), Claude 3 Sonnet and Opus showed notably larger distances (1.2). The Gemini family similarly exhibited significant version-dependent variations, with Gemini 1.5 Preview (April 2024) showing the minimum distance of 0.9, while subsequent versions sometimes demonstrated larger distances of 1.1. Additionally, 1.5 Flash models showed greater distances compared to 1.5 Pro models.
Among open-source models, the Gemma family’s latest Gemma 2 2B and 9B models showed notably large distances from human judgments (2.0), with distances decreasing as model size increased. The Llama family demonstrated similar trends, with 70B models showing closer alignment to human judgments compared to smaller (1–8B) models.
Regarding other LLMs, Mistral-Nemo and Command R+ demonstrated relatively close alignment with human judgments (distances 0.8–0.9), while Vicuna and Phi 3.5 showed larger distances (1.0–1.6).
Factors influencing judgment distance
Proprietary vs. open-source models
To identify factors influencing the distance from human judgments, we first examined differences based on model types. This investigation was motivated by the general assumption that proprietary models demonstrate superior performance compared to open-source models [50], suggesting potential differences in their alignment with human judgments.
Models were classified into three groups: proprietary models, open-source models, and open-source models with parameters exceeding 10B (Fig 3). Proprietary models showed a median distance of 0.9 from human judgments, while open-source models as a whole demonstrated larger distances with a median of 1.2 (Wilcoxon test’s $p$-value $p=0.0081$).
However, to determine whether this difference was fundamentally attributable to the proprietary/open-source distinction or merely reflected differences in model sizes, we conducted additional analysis focusing on open-source models with parameters exceeding 10B. Results showed that large-scale open-source models exhibited a median distance of 0.9 from human judgments, with no significant difference from proprietary models (Wilcoxon test’s $p=0.93$).
These findings suggest that differences in judgment distance from humans depend more strongly on model size rather than the proprietary/open-source distinction.
Fig 3: Comparison of moral judgment distances between proprietary and open-source models Violin plots with embedded box plots compare the distribution of distances from human judgments across three model categories: proprietary models, all open-source models, and large open-source models with parameters exceeding 10B. Individual models are represented as points.
Impact of model size
We conducted a detailed analysis of the relationship between model size and distance from human judgments, focusing on open-source models (Fig 4). Results revealed a significant negative correlation between model size and distance from human judgments (Spearman’s rank correlation coefficient $\rho=-0.50$, $p=0.018$), indicating that larger models tend to align more closely with human judgments.
Particularly in the Llama family, while 1-3B scale models (Llama 3.2 1B, 3B) showed distances of 1.2-1.5, 70B scale models (Llama 3 70B, 3.1 70B) demonstrated notably smaller distances of 0.7. Similar trends were observed in the Gemma family, where 2-9B scale models from Gemma 1.1 onward showed relatively large distances of 1.2-1.9, while the Gemma 2 27B scale model demonstrated a smaller distance of 1.1.
However, as indicated by the moderate correlation coefficient, the relationship between model size and distance from human judgments is not straightforward. For instance, within the Llama family, while Llama 3.2 1B showed a distance of 1.2, Llama 3.2 3B demonstrated a distance of 1.5. Similarly, among 8B models, version 3 showed a distance of 0.8, while version 3.1 exhibited a distance of 1.6. In the Gemma family, while Gemma 2B showed the minimum distance, subsequent larger models demonstrated increased distances. These observations suggest that while model size is a significant factor influencing judgment distance from humans, other factors (such as model architecture and training methodology) may play equally important roles.
Note that proprietary models were excluded from this analysis due to limited information about their sizes. However, given that proprietary models are generally believed to have parameters exceeding several hundred billion and demonstrate small median distances of 0.9 from human judgments as previously mentioned, this correlation would likely persist even if proprietary models were included.
Fig 4: Relationship between model size and distance from human judgments in open-source models. Different model families are represented by different colors. Model names are labeled. Horizontal axis (model size in billion parameters) are represented in logarithmic scale.
Impact of model updates
While previous analyses suggested model size as a significant factor influencing distance from human judgments, our observations indicated that updates within the same model family do not necessarily reduce judgment distances. Therefore, we conducted a more detailed analysis of how model updates affect distances from human judgments.
The impact of updates on open-source LLMs can be observed in detail from Fig 4. For instance, examining 2B Gemma models revealed increasing distances from human judgments through version progression from Gemma 2B to 1.1 2B and 2 2B. Similar trends were observed in 7–9B Gemma models, suggesting that updates do not necessarily lead to closer alignment with human judgment patterns.
The Llama family also demonstrated non-uniform effects of updates. Among 7–8B models, while the update from Llama 2 to 3 decreased distance from 1.1 to 0.8, showing closer alignment with human judgments, the subsequent update to Llama 3.1 increased distance to 1.6. Conversely, larger 70B models maintained a consistent distance of 0.7 through the update from Llama 3 to 3.1.
For proprietary models, we conducted temporal analysis of the relationship between release dates and distances from human judgments (Fig 5), focusing on GPT-3.5, GPT-4, GPT-4o, Gemini Pro, Gemini Flash, and Claude Sonnet families, which underwent multiple updates.
The GPT-3.5 and GPT-4 families maintained distances between 0.6–0.9 from March 2023 to April 2024. GPT-4 achieved its minimum distance of 0.6 with the January 2024 version but increased to 0.8 in the April version. GPT-3.5 showed its minimum distance of 0.7 in the June 2023 version, increased slightly to 0.9 in the November 2023 version, then returned to 0.8 in January 2024.
The newer GPT-4o family (from May 2024) showed slightly larger distances of 0.9–1.0 compared to traditional GPT-4. The Gemini family maintained relatively small distances of 0.9 through 1.0 Pro to 1.5 Pro Preview (April 2024), but later versions increased to 1.1. Gemini 1.5 Flash showed large initial distances of 1.2 (Pre, 001 versions) but decreased to 0.9 in version 002. The Claude Sonnet family’s distance decreased from 1.2 to 0.9 with the update from 3 to 3.5 (June 2024) but increased to 1.0 with the October 2024 update of 3.5. The Claude Haiku family’s distance increased from 0.9 to 1.1 with the update from 3 to 3.5.
These observations indicate that updates, whether in open-source or proprietary models, do not necessarily lead to closer alignment with human judgments. Furthermore, newer versions or supposedly higher-performing models do not consistently demonstrate closer alignment with human judgments.
Fig 5: Temporal changes in distance from human judgments across proprietary model families. Different model families are represented by different colors. Model names are labeled.
Comprehensive analysis of moral judgments
We conducted PCA using AMCE values across nine categories as feature vectors to comprehensively evaluate similarities in judgment tendencies among models (Fig 6). The first and second principal components explained 74.2% of total variance (46.1% and 28.1% respectively), enabling visualization of judgment similarities in this two-dimensional plane.
This analysis revealed several characteristic patterns. First, the GPT-4 family, GPT-4o/o1 family, Claude 3 Haiku, Claude 3.5 Sonnet, Gemini Pro 1.5 family, and GPT-3.5 family formed clusters relatively close to human judgments. Notably, these models concentrated in the positive region of the first principal component, suggesting commonalities in their judgment patterns.
Conversely, Claude 3 Sonnet and Opus, along with Gemini 1.5 Flash Pre and Flash 001, positioned in strongly negative regions of the first principal component, demonstrating judgment patterns distinct from humans and other proprietary models. The Gemma 2 family (2B, 9B) showed large positive values in the second principal component, confirming distinctively different judgment patterns from other models.
Interestingly, these results aligned with our previous distance analysis in multiple aspects. Models showing smaller distances from human judgments tended to position closer to humans in this two-dimensional plane. Similar trends were observed regarding model size effects. For instance, within the Llama family, 70B models (Llama 3 70B, 3.1 70B) positioned near humans, while smaller 1-8B models distributed farther away.
Changes due to model updates were also clearly observable in this two-dimensional plane. For example, GPT-3.5 approached human positions most closely in the June 2023 version but moved away with subsequent updates. Similarly, while Claude Sonnet moved closer to humans with the update from 3 to 3.5, it slightly diverged with the October version of 3.5. These trajectories support our previous finding that updates do not necessarily lead to closer alignment with human judgments.
Fig 6: Principal component analysis of LLM moral judgments compared to human preferences. the first two principal components derived from AMCE values across nine moral preference categories. Different model families are represented by different colors, and human judgment is marked with a black diamond.
Discussion
In this study, we significantly expanded upon the four LLMs analyzed in our previous study [33] to comprehensively examine ethical judgments in autonomous driving across 51 different LLMs using the Moral Machine experimental framework (Figs 1 and 2). Proprietary models and open-source models exceeding 10B parameters demonstrated relatively small distances from human judgments (Figs 3 and 4), showing consistent alignment with human preferences in fundamental moral decisions, such as prioritizing humans over animals and favoring the preservation of more lives. However, model updates did not necessarily lead to closer alignment with human judgments, and variations in judgment tendencies were observed across updates (Fig 5). This extensive model comparison enabled a more comprehensive understanding of the evolution of moral judgments in LLMs.
The incomplete alignment between LLM and human moral judgments is an expected outcome, considering that these models are not explicitly trained to mimic human moral judgments. However, the significant negative correlation observed between model size and distance from human judgments in open-source models (Fig 4) warrants attention. This finding aligns with scaling laws widely observed in language model behavior, suggesting that merely increasing model size might enable LLMs to naturally acquire human-like moral judgments. Nevertheless, this correlation does not necessarily imply causation, and other factors such as differences in model architecture, training data quality, and optimization methods must be considered. For instance, we observed cases within the same model family where size increases did not result in monotonic decreases in distance from human judgments.
Differences in judgment tendencies across model families provide important implications for implementing autonomous driving systems. While the GPT-4 family demonstrated the closest alignment with human judgments (distances 0.6–0.8), its proprietary nature and black-box internal structure present implementation challenges. Among open-source models, Llama 70B showed relatively close alignment with human judgments (distances 0.7–0.8), making it promising from the perspective of balancing model transparency and ethical judgment quality. However, autonomous driving systems often require real-time decision-making, and computational cost and latency constraints make lighter models more practical. In this regard, smaller models face the challenge of relatively larger distances from human judgments (1.2–1.5), indicating a clear trade-off between judgment quality and practicality.
Addressing this trade-off remains a crucial challenge for future development. While several LLMs fine-tuned specifically for autonomous driving tasks have been proposed [17], showing progress in technical aspects such as perception, control, and navigation, their ethical judgment capabilities and moral reasoning have not been sufficiently evaluated. Implementing small-scale models in practical systems requires developing fine-tuning methods that suppress excessive emphasis on specific judgment criteria and establishing adaptive learning methods that consider cultural backgrounds. Furthermore, monitoring mechanisms for regularly evaluating and adjusting model judgment tendencies will be crucial for real-world operations.
Divergences between LLM and human judgments primarily manifested as excessive emphasis on ethical principles. While many LLMs aligned with humans in prioritizing humans over pets and pedestrians over passengers, the magnitude of these preferences substantially exceeded human values. These tendencies align with the German Ethics Code on Automated and Connected Driving’s [51] Ethical Guideline 7, particularly regarding prioritizing the prevention of personal injury over damage to animals and property, and Ethical Guideline 5, particularly regarding defensive driving for vulnerable road users, though their excessive application raises new concerns.
Particularly notable is the extreme emphasis on maximizing lives saved. This can be interpreted as a reflection of Western individualistic and utilitarian values [10, 52], suggesting that LLMs may over-emphasize these cultural characteristics, likely due to training data predominantly sourced from Western cultures [53, 54]. However, the German Ethics Code’s Ethical Guideline 2 explicitly rejects purely utilitarian approaches in autonomous vehicle decision-making, advocating against judgments based solely on maximizing the number of lives saved.
LLM judgment tendencies demonstrated both alignment and divergence from human preferences. Preferences favoring overweight over physically fit individuals and lower over higher social status individuals showed inverse patterns to human judgments, potentially interpreted as over-application of principles protecting vulnerable individuals. Conversely, while preferences for females over males and law-abiding over law-breaking individuals aligned directionally with human judgments, their magnitude significantly exceeded human values. These can be understood as reflections of the historical “Women and Children First” principle [55] and principles of fairness and justice [56], respectively. These judgment patterns suggest that LLMs extract specific social values and ethical principles from training data and apply them in extreme forms.
These findings raise significant challenges for implementing LLMs in autonomous driving systems. While human judgments demonstrate flexibility and context-dependency, LLM judgments appear rigid.
These findings raise significant challenges for implementing LLMs in autonomous driving systems. While human judgments demonstrate flexibility and context-dependency, LLM judgments appear rigid. Recent research [57] has further highlighted this concern by demonstrating that LLMs exhibit different moral biases across languages, suggesting that their ethical decision-making is not consistently aligned even within the same model. This linguistic variability in moral judgments raises additional concerns about the global deployment of LLM-based autonomous systems.
Considering that AI system decisions may influence human decision-making[30], the use of LLMs could potentially amplify specific moral perspectives and exacerbate social divisions. This risk is particularly acute given that LLMs’ moral preferences can vary significantly across languages and cultures. Moreover, these judgment tendencies fundamentally conflict with the German Ethics Code’s Ethical Guideline 9 of equal treatment regardless of demographic factors and identity.
The excessive reflection of Western values and extreme application of specific ethical principles may reduce the acceptability of autonomous driving systems in a culturally diverse global society. This concern is particularly relevant given the evidence that LLMs’ moral judgments can diverge significantly from human preferences across different cultural and linguistic contexts. Addressing these challenges requires appropriate control of LLM judgment tendencies and the design of more flexible decision-making mechanisms that consider cultural and linguistic diversity.
This study has several important limitations. In evaluating LLM judgment tendencies, we used globally aggregated moral preferences as the human judgment benchmark. However, inherently moral choices about whom to save are known to be significantly influenced by cultural and social factors. For instance, utilitarian judgment criteria favoring the preservation of more lives may have varying levels of acceptance across different cultural spheres. Our analysis did not conduct detailed examinations of such cultural and social differences. When integrating AI into autonomous driving systems, evaluation must not only align AI judgment tendencies with human values but also consider regional cultural and social contexts.
The Moral Machine experiment itself, based on trolley-type dilemmas, also has limitations. As discussed in [33], this experimental methodology faces issues of real-world disconnection and binary choice simplification. Additionally, it cannot incorporate judgments based on road users’ personality and intentions, nor can it evaluate low-risk moral judgments in everyday traffic situations. For greater ecological validity, new experimental paradigms have been proposed combining virtual reality-based everyday traffic moral scenarios with moral psychological frameworks such as the Agent-Deed-Consequences (ADC) model [58].
In this context, multimodal large language models (MLLMs) offer a promising approach to realizing such new experimental paradigms. Recently, MLLMs have shown potential for application in autonomous driving systems, integrating visual information and language understanding for situation awareness and decision-making [18, 19]. While conventional LLMs make judgments based solely on textual descriptions of situations, MLLMs can directly process visual information from traffic scenes. This capability enables the evaluation of ethical judgments in realistic traffic scenarios created through virtual reality, facilitating more comprehensive analysis.
Despite these limitations, our comprehensive analysis of 51 different LLMs provides important insights into the characteristics of LLM ethical judgments in autonomous driving systems. Specifically, our findings on the relationship between model size and ethical judgment, characterization of judgment tendencies across model families, and identification of concrete implementation challenges will likely serve as crucial guidelines for the ethical design of future autonomous driving systems. Furthermore, our perspective on more realistic evaluation methods, acknowledging the constraints of existing experimental paradigms, is expected to contribute to further development in this field.
Supporting information
S1 Fig.
Differences in AMCE values between LLMs and human preferences. Heatmap showing the differences between LLM and human AMCE values for nine moral preference categories.
S1 Table.
AMCE values for nine moral preference categories across all analyzed LLMs.
S2 Table.
Comprehensive model information and evaluation metrics. Summary of all analyzed LLMs, including their characteristics and performance metrics. The ‘open-source’ column indicates model accessibility (yes/no), ‘Size’ shows the number of parameters in billions where available, ‘Date’ indicates the model version’s release date, ‘Distance’ shows the Euclidean distance between the model’s and human AMCE values across nine moral preference categories, and ‘valid response rate’ represents the proportion of valid responses in the evaluation scenarios.

### v2

@@
 While our previous study examined four prominent LLMs using the Moral Machine experimental framework, the dynamic landscape of LLM development demands a more comprehensive analysis.
-Here, we evaluate moral judgments across 51 different LLMs, including multiple versions of proprietary models (GPT, Claude, Gemini) and open-source alternatives (Llama, Gemma), to assess their alignment with human moral preferences in autonomous driving scenarios.
+Here, we evaluate moral judgments across 52 different LLMs, including multiple versions of proprietary models (GPT, Claude, Gemini) and open-source alternatives (Llama, Gemma), to assess their alignment with human moral preferences in autonomous driving scenarios.
 Using a conjoint analysis framework, we evaluated how closely LLM responses aligned with human preferences in ethical dilemmas and examined the effects of model size, updates, and architecture.
@@
 Introduction
-The rapid advancement of Large Language Models (LLMs) has inaugurated a new era in artificial intelligence (AI), demonstrating unprecedented capabilities across diverse domains [1, 2, 3].
+The rapid advancement of Large Language Models (LLMs) has inaugurated a new era in artificial intelligence (AI), demonstrating unprecedented capabilities across diverse domains [1, 2, 3, 4].
 These models, exemplified by ChatGPT [5, 6] (developed by OpenAI) and its counterparts, exhibit remarkable proficiency in natural language processing, content generation, complex problem-solving, and decision-making.
-As LLMs evolve, their potential integration into critical systems has become a subject of intense scrutiny and research [6, 7, 8].
-Autonomous driving represents a field where LLM application holds particular promise while raising significant ethical questions [9, 10, 11, 12].
-The automotive industry is actively exploring LLM integration to enhance autonomous vehicles’ capabilities, potentially improving environmental comprehension, complex scenario interpretation, nuanced decision-making, and communication with passengers and other road users [13, 14, 15, 16, 17, 18, 19].
-This integration spans various aspects of autonomous driving, including perception, navigation, planning, and control.
-The Moral Machine experiment [10] emphasized the importance of ethical considerations in AI decision-making for autonomous systems.
-This study presented participants with scenarios involving unavoidable accidents to assess public opinion on how autonomous vehicles should act in morally challenging situations.
-The experiment provided a framework for exploring human moral preferences, identifying trends such as prioritizing human lives over animals and favoring the preservation of a greater number of lives.
-While acknowledging its limitations [20, 21, 22, 23, 24, 25], this work has stimulated extensive research into the ethical implications of AI in autonomous systems [26, 27, 28, 29].
+As these models advance from laboratory experiments to real-world applications, understanding their decision-making processes, particularly in ethical contexts, becomes crucial for ensuring their responsible deployment.
+This has become a subject of intense scientific investigation [7, 8, 9].
+Autonomous driving represents a field where LLM application necessitates particularly careful evaluation [10, 11, 12, 13].
+When autonomous vehicles make decisions on public roads, their choices directly affect human safety and well-being.
+The automotive industry’s investigation of LLM integration for enhancing autonomous vehicles’ capabilities includes various aspects: analyzing complex road conditions, interpreting traffic situations, planning appropriate responses, and importantly, making ethical decisions in challenging scenarios [14, 15, 16, 17, 18, 19, 20].
+This integration spans critical functions from environmental comprehension to decision making, highlighting the need for thorough understanding of these models’ capabilities and limitations.
+The Moral Machine experiment [11] provides a systematic framework for evaluating ethical decision-making in autonomous systems.
+This study presented participants with scenarios involving unavoidable accidents to assess how autonomous vehicles should behave in morally challenging situations.
+The experiment revealed clear patterns in human moral preferences, such as prioritizing human lives over animals and favoring the preservation of a greater number of lives.
+While acknowledging its limitations including methodological constraints in scenario design, challenges in cultural representation, and gaps between theoretical choices and practical implementation [21, 22, 23, 24, 25, 26], this work has stimulated extensive research into the ethical implications of AI in autonomous systems [27, 28, 29, 30].
+These foundational insights, despite their limitations which we explore in detail in our discussion, have provided crucial frameworks for examining ethical decision making in autonomous systems.
 While some studies have explored AI responses to standard ethical dilemmas [31, 32], such as the classic trolley problem [33], the intricate scenarios presented in the Moral Machine experiment offer a more comprehensive exploration of AI moral preferences.
@@
 Moral Machine scenario generation
-Following the methodology established in our previous study [33], we generated a comprehensive set of Moral Machine scenarios to evaluate LLM ethical decision-making.
+To comprehensively evaluate how different LLMs approach moral decision-making in autonomous driving scenarios and assess the effects of model updates on these judgments, we generated a systematic set of test scenarios following our previous methodology.
 These scenarios presented dilemmas involving an autonomous vehicle facing an unavoidable accident, requiring a choice between two outcomes (“Case 1” and “Case 2”) with varying ethical implications.
@@
 Large language models
-This study expanded upon our previous study [33] by incorporating a comprehensive set of large language models (LLMs), including the latest versions available.
+To investigate how moral judgment patterns vary across different model architectures and evolve through updates, we expanded upon our previous study [34] by incorporating a comprehensive set of LLMs, including the latest versions available.
+Our analysis encompasses both base language models and models specifically optimized for dialogue interactions.
+While these models may differ in their training approaches, they share fundamental capabilities in processing and responding to complex queries, including moral scenarios.
 These LLMs were utilized to generate responses to moral machine scenarios.
@@
 From Anthropic, we evaluated the latest versions of their Claude models, including Claude 3 Haiku, Sonnet, Opus [35], and Claude 3.5 Sonnet [42].
-In the open-source domain, we included Meta’s Llama family [36], comprising Llama 3 and Llama 3.1 in 8B and 70B versions, as well as Llama 3.2 in 1B and 3B versions.
+In the open-source domain, we included Meta’s Llama family [37], comprising Llama 3 and Llama 3.1 in 8B and 70B versions, Llama 3.2 in 1B and 3B versions, and Llama 3.3 in 70B version.
 A comprehensive evaluation of Google’s Gemma model family [38, 43] was conducted, encompassing multiple versions and scales from 2B-it to 27B-it, including the derivative DataGemma RIG 27B-it model [44].
@@
 For a comprehensive list of all models and their specific versions used in this study, please refer to our GitHub repository (https://github.com/kztakemoto/mmllm).
-All models were used with their default parameter settings. The input prompts were consistent with those used in [33], requiring each model to select either “Case 1” or “Case 2” for each scenario.
+All models were used with their default parameter settings.
+The input prompts were consistent with those used in [34], requiring each model to select either “Case 1” or “Case 2” for each scenario.
+The complete code for prompt generation, analysis, and resulting data is available in our GitHub repository.
 While most models were evaluated using 50,000 scenarios, following our previous study [34], constraints related to application programming interface (API) usage costs and computational time necessitated the use of fewer scenarios for certain models.
@@
 Data analysis
-For analyzing LLM responses to the Moral Machine scenarios, we employed the conjoint analysis framework as described in [49] and implemented in the original study [10] on the Moral Machine experiment and our previous study [33].
+To systematically evaluate how different LLMs align with human moral preferences and how this alignment changes across model versions and architectures, we analyzed the LLM responses using the conjoint analysis framework as described in [50] and implemented in the original study [11] on the Moral Machine experiment and our previous study [34].
 This approach allows for a robust, non-parametric identification of causal effects without specific modeling assumptions.
@@
 Moral judgments of large language models
-Figure 1 presents radar charts characterizing the moral judgment tendencies of various LLMs across nine preferences quantified by AMCE values (see S1 Table.
+Moral judgments of large language models To comprehensively evaluate how different LLMs approach moral decision-making in autonomous driving scenarios, we conducted a systematic analysis of their responses across multiple ethical dimensions.
+Our analysis framework focused on quantifying and comparing the moral preferences of various LLM families with human judgments.
+Fig 1 presents radar charts characterizing the moral judgment tendencies of various LLMs across nine preferences quantified by Average Marginal Component Effect (AMCE) values (see S1 Table.
 for exact numerical data).
-These charts compare the preferences across multiple versions and model sizes of GPT-3.5, GPT-4, GPT-4o/o1, Claude, Gemini, Llama, Gemma, and other LLMs against human preferences (see S1 Fig.
+These charts enable direct comparison of preferences across multiple versions and model sizes of major LLM families, including GPT-3.5, GPT-4, GPT-4o/o1, Claude, Gemini, Llama, Gemma, and other LLMs, against established human preferences (see S1 Fig.
 for detailed differences between each model and human values).
-Analysis revealed that many LLMs demonstrated large positive values (¿0.5) for No.
-Characters and Species dimensions, indicating tendencies similar to human judgments in prioritizing the preservation of more lives and favoring humans over animals.
-However, some models exhibited either inverse priorities compared to humans or disproportionately strong preferences in the same direction (differences in AMCE values approximately ¿0.3 compared to human values).
-We elaborate on these findings below, including the effects of model updates.
+Our initial analysis revealed three key patterns in LLM moral judgments.
+First, many LLMs demonstrated strong alignment with human preferences in fundamental moral decisions, showing large positive values ($>0.5$) for No.
+Characters and Species dimensions.
+This indicates consistent tendencies in prioritizing the preservation of more lives and favoring humans over animals, suggesting that LLMs can capture basic human moral intuitions.
+Second, we observed significant variations in judgment patterns across different model families, with some models exhibiting inverse priorities compared to human preferences.
+Third, certain models showed disproportionately strong preferences in the same direction as humans (differences in AMCE values approximately $>0.3$ compared to human values), indicating potential overemphasis of specific moral principles.
+These patterns provide important insights into both the capabilities and limitations of LLMs in moral reasoning tasks.
+Below, we elaborate on these findings for each model family, examining how different architectures influence moral judgments, and analyzing the effects of model updates on alignment with human preferences.
 {adjustwidth}
@@
 GPT-3.5 family
-Analysis of the GPT-3.5 family (Fig 1A) revealed interesting changes associated with model updates.
-While models from June 2023 to January 2024 exhibited similar preference patterns, the March 2023 version demonstrated distinctive judgment tendencies.
-For instance, regarding Law, the March version showed positive values favoring law-abiding individuals, aligning with human preferences but displaying stronger preferences than humans.
-However, subsequent versions shifted toward more neutral judgments with values approaching zero.
-Similar changes were observed in Gender-related judgments.
-The March 2023 version showed strong positive values favoring females, but subsequent updates gradually reduced these values, converging toward human judgment levels.
-Regarding age-related decisions, while humans demonstrated a preference for younger individuals, all GPT-3.5 versions maintained neutral judgments with values near zero.
-For Fitness-related judgments, the March and June 2023 versions showed negative values contrary to human preferences, prioritizing overweight individuals over physically fit ones.
-However, later versions gradually aligned more closely with human judgments.
-Regarding Relation to AV, all versions consistently showed positive values similar to humans, but with notably higher magnitudes, indicating a stronger preference for pedestrians over passengers.
-For Species-related judgments, while the March and June 2023 versions showed values comparable to humans, the November 2023 and January 2024 versions exhibited decreased values, maintaining but weakening the tendency to prioritize humans over pets.
+Analysis of the GPT-3.5 family (Fig 1A) revealed significant temporal evolution in moral judgments across model updates.
+While models from June 2023 to January 2024 exhibited similar patterns, the March 2023 version showed distinctive tendencies, particularly in several key ethical dimensions.
+In the Law dimension, the March version showed strong positive values favoring law-abiding individuals (AMCE = 0.7), but subsequent versions shifted toward more neutral judgments (AMCE $\approx$ 0).
+Similar temporal changes were observed in Gender-related judgments, where initial strong female preference (AMCE = 0.5 in March 2023) gradually reduced toward human judgment levels (AMCE = 0.2 in January 2024, human AMCE = 0.1).
+Fitness-related judgments also evolved from initially contradicting human preferences (AMCE $=-0.2$ in March 2023) to showing better alignment in later versions (AMCE = 0.1 in January 2024).
+However, we observed persistent deviations from human preferences in certain dimensions.
+The latest versions maintained near-neutral age-related judgments (AMCE $\approx 0$) despite human preference for younger individuals (AMCE = 0.5), and showed consistently stronger preferences for pedestrians over passengers (AMCE = 0.6 vs.
+human 0.1).
+Species-related judgments demonstrated an interesting pattern where initial alignment with human values in earlier versions (AMCE = 0.7) weakened in later updates (AMCE = 0.2) while maintaining the same directional preference.
 GPT-4 family
-Analysis of the GPT-4 family (Fig 1B) revealed high consistency across models from March 2023 to January 2024, demonstrating that model updates had limited impact on moral judgments.
-Across all nine preference categories analyzed, the models exhibited directional alignment with human judgments, though showing stronger preferences in specific categories.
-Most notably, in the Species and No.
-Characters categories, AMCE values approached 1.0, indicating nearly deterministic preferences for saving humans over animals and larger groups over smaller ones.
-Similarly, in the Gender, Law, and Relation to AV categories, while aligned directionally with human judgments, the models displayed larger AMCE values.
-Specifically, they showed more pronounced tendencies than humans in prioritizing females over males, law-abiding individuals over lawbreakers, and pedestrians over passengers.
-Conversely, in the Fitness category, the models demonstrated trends diverging from human judgments.
-While humans showed a slight preference for physically fit individuals, the GPT-4 family exhibited a minor inverse preference, marginally favoring overweight individuals.
-However, this divergence was relatively minor compared to the pronounced preferences observed in other categories.
+Analysis of the GPT-4 family (Fig 1B) revealed remarkable consistency in moral judgments across different versions from March 2023 to January 2024.
+In contrast to GPT-3.5, model updates showed minimal impact on judgment patterns, with all versions maintaining similar preference strengths across ethical dimensions.
+The GPT-4 family demonstrated strong alignment with human judgments in direction but often exceeded human preference magnitudes.
+Most notably, Species and No.
+Characters preferences showed near-deterministic values (AMCE $>0.9$) compared to human values (AMCE = 0.6), indicating substantially stronger preferences for saving humans over animals and larger groups over smaller ones.
+Similar amplification appeared in Gender (AMCE = 0.4 vs.
+human 0.1), Law (AMCE = 0.6 vs.
+human 0.3), and Relation to AV (AMCE = 0.5 vs.
+human 0.1) categories.
+The only notable divergence appeared in the Fitness category, where GPT-4 models showed a slight preference for overweight individuals (AMCE $=-0.1$) contrary to human preference for physically fit individuals (AMCE = 0.2).
+This deviation, however, was minimal compared to the strong alignment observed in other dimensions.
 GPT-4o/o1 family
-Analysis of the GPT-4o family (Fig 1C) similarly demonstrated high consistency across models, with model updates showing limited impact on moral judgments.
-While this family exhibited AMCE values comparable to human judgments in the Gender and No.
-Characters categories, notable differences from human preferences were observed across many categories.
-Particularly in the Relation to AV category, all models except GPT-4o-mini showed substantially higher values compared to even the GPT-4 family, demonstrating an extreme preference for pedestrians over passengers.
-The strong preference for law-abiding individuals persisted, and all GPT-4o models showed significantly stronger tendencies than humans in prioritizing humans over pets.
-Conversely, these models demonstrated relatively weak preferences for younger individuals compared to human age-related judgments, while maintaining a slight, albeit weak, preference for overweight individuals in Fitness-related decisions.
-The o1 model generally exhibited judgment patterns similar to the GPT-4o family, though the o1-mini model showed distinctive variations.
-Notably, o1-mini demonstrated near-neutral preferences aligned with human judgments in the Relation to AV category.
-However, it recorded AMCE values approaching 1.0 in the No.
-Characters category, indicating a strong preference for saving larger groups of lives.
+Analysis of the GPT-4o family (Fig 11) revealed a pattern of consistent moral judgments across models, with two notable exceptions: GPT-4o-mini and o1-mini showed distinctive characteristics.
+The family exhibited closest alignment with human preferences in Gender (AMCE = 0.1) and No.
+Characters (AMCE = 0.7) categories, matching human values within ±$\pm 0.1$ AMCE.
+Most GPT-4o models showed stronger ethical preferences than both humans and the GPT-4 family in several dimensions.
+In Relation to AV, all models except GPT-4o-mini demonstrated extreme pedestrian preference (AMCE $>0.5$ vs.
+human 0.1).
+Similarly, they showed amplified preferences for law-abiding individuals (AMCE = 0.8 vs.
+human 0.4) and humans over pets (AMCE = 0.9 vs.
+human 0.6).
+However, these models showed weaker age-related preferences (AMCE = 0.2 vs.
+human 0.5) and maintained slight preferences for overweight individuals in Fitness judgments (AMCE $=-0.1$).
+The o1 variants showed interesting divergences: while the standard o1 followed GPT-4o patterns, o1-mini demonstrated unique characteristics, including near-neutral Relation to AV preferences (AMCE = 0.1) but extremely strong preferences for larger groups (AMCE $\approx$ 1.0).
+This suggests that model size significantly influences moral judgment patterns within this family.
 Claude family
-Analysis of the Claude family (Fig 1D) revealed a distinct bifurcation in judgment tendencies across models.
-Specifically, two characteristic groups emerged: one comprising Claude 3 Sonnet, Opus, and 3.5 Haiku, and another consisting of Claude 3 Haiku and Claude 3.5 Sonnet.
-Claude 3 Sonnet, Opus, and 3.5 Haiku demonstrated distinctive trends that diverged from human judgments across multiple categories.
-These models showed negative AMCE values in Relation to AV, Fitness, and Social Status categories, indicating inverse preferences compared to humans: prioritizing passengers over pedestrians, overweight over physically fit individuals, and lower over higher social status.
-While these models aligned directionally with humans in Gender (favoring females over males) and Species (prioritizing humans over pets) categories, the magnitude of their preferences substantially exceeded human values.
-Conversely, Claude 3 Haiku and Claude 3.5 Sonnet exhibited different patterns, including neutral age-related judgments that diverged from human preferences.
-These models also demonstrated excessive preferences, particularly in Relation to AV and Law categories, showing substantially higher AMCE values than humans in prioritizing pedestrians over passengers and law-abiding over law-breaking individuals.
-Claude 3.5 Sonnet notably demonstrated pronounced tendencies in Species and No.
-Characters categories, strongly favoring humans over pets and larger over smaller groups.
-An interesting finding regarding model evolution revealed that the Sonnet model’s update from 3 to 3.5 generally resulted in closer alignment with human judgments.
-However, no significant changes in judgment patterns were observed between the June and October versions of Claude 3.5 Sonnet.
-Conversely, the Haiku model’s update from 3 to 3.5 resulted in greater divergence from human judgments.
+Analysis of the Claude family (Fig 1D) revealed a clear bifurcation in moral judgment patterns, forming two distinct groups with contrasting ethical preferences.
+The first group (Claude 3 Sonnet, Opus, and 3.5 Haiku) and second group (Claude 3 Haiku and Claude 3.5 Sonnet) showed markedly different alignment with human preferences.
+The first group demonstrated significant divergence from human judgments in several dimensions.
+They showed inverse preferences in Relation to AV (AMCE $=-0.4$), Fitness (AMCE $\approx-0.5$), and Social Status (AMCE $=-0.3$) categories, contrasting with human values (AMCE = 0.1, 0.2, and 0.3 respectively).
+While they aligned with humans in Gender and Species preferences, their magnitude was notably stronger in both dimensions (Gender: AMCE = 0.7 vs.
+human 0.3; Species: AMCE = 0.9 vs.
+human 0.6).
+The second group showed different patterns: Claude 3 Haiku and Claude 3.5 Sonnet exhibited excessive preferences in Relation to AV (AMCE = 0.5) and Law (AMCE = 0.7) categories compared to humans (AMCE = 0.1 and 0.4, respectively).
+Claude 3.5 Sonnet particularly showed strong preferences in Species and No.
+Characters dimensions (AMCE $>0.8$).
+Model updates showed inconsistent effects: while the Sonnet 3 to 3.5 update improved human alignment, the Haiku 3 to 3.5 update increased divergence.
+Subsequent updates within Claude 3.5 Sonnet showed minimal changes in judgment patterns.
 Gemini family
-Analysis of the Gemini family (Fig 1E) revealed high inter-model consistency across most categories, while demonstrating notable changes in judgment tendencies with updates, particularly in the Relation to AV category.
-Specifically, while the Gemini 1.0 Pro model showed negative AMCE values indicating preference for passengers over pedestrians, the Gemini 1.5 Pro family shifted to positive values aligned with human judgments, prioritizing pedestrians over passengers.
-Similar update-driven changes were observed in the Gemini 1.5 Flash family.
-While the 1.5 Flash Preview and 1.5 Flash 001 versions prioritized passengers, the update to 1.5 Flash 002 reversed this preference to favor pedestrians.
-However, the degree of pedestrian preference in the latest versions of both Gemini 1.5 Pro and 1.5 Flash substantially exceeded human values.
-Regarding gender-related judgments, only the latest Gemini 1.5 Pro 002 showed values comparable to human judgments, while other versions demonstrated pronounced preferences for females over males.
-In the Species category, all versions showed stronger preferences than humans for prioritizing humans over pets.
-Conversely, all models consistently demonstrated inverse tendencies to humans in Fitness-related decisions, showing negative AMCE values favoring overweight over physically fit individuals.
-Figure 1E includes PaLM 2, Google’s previous proprietary model, for comparison.
-PaLM 2 exhibited a distinctive tendency contrary to human judgments, prioritizing smaller over larger groups.
-In contrast, the Gemini family, Google’s current flagship models, aligned with human judgments across all versions in prioritizing the preservation of more lives.
+Analysis of the Gemini family (Fig 1E) revealed significant evolution in moral judgments across model updates, particularly marked by changes in the Relation to AV dimension.
+We observed a clear transition from Gemini 1.0 Pro to the 1.5 series, with distinct changes in ethical preferences.
+The most notable shift occurred in the Relation to AV category, where Gemini 1.0 Pro initially showed preference for passengers (AMCE $=-0.4$), while the Gemini 1.5 Pro series shifted to strongly favor pedestrians (AMCE = 0.7 vs.
+human 0.1).
+The Gemini 1.5 Flash family showed a similar evolution: Flash Preview and 001 favored passengers (AMCE $=-0.7$), but Flash 002 shifted to strong pedestrian preference (AMCE = 0.7).
+Gender and Species preferences also showed distinctive patterns.
+Only Gemini 1.5 Pro 002 aligned closely with human gender preferences (AMCE = 0.2 vs.
+human 0.1), while earlier versions showed stronger female preference (AMCE $>0.4$).
+All versions showed strong preferences than humans for prioritizing humans over pets (AMCE $>0.7$) and consistently showed inverse fitness-related preferences (AMCE $=-0.3$ vs.
+human 0.2).
+Notably, compared to Google’s previous PaLM 2 model, which contradicted human preferences in group size decisions (AMCE = $-0.3$ vs.
+human 0.7), all Gemini versions maintained consistent alignment with human preferences in preserving larger groups (AMCE = 0.4–0.8).
 Llama family
-Analysis of the Llama family (Fig 1F) demonstrated that all models except Llama 3.1 8B showed positive AMCE values similar to humans in the No.
-Characters category, indicating consistent preferences for saving larger groups of lives.
-However, other categories exhibited varying judgment patterns across versions and model sizes.
-For gender-related decisions, while Llama 2 (7B), 3.2 1B, and 3.2 3B models showed preferences for males contrary to human judgments, other models aligned with humans in favoring females.
-In the Species category, Llama 3 70B, 3.1 8B, and 3.1 70B models demonstrated clear preferences for humans over pets aligned with human judgments, while other models showed either neutral judgments or slight preferences for pets.
-Similarly, regarding law compliance, Llama 3 8B, 3 70B, and 3.1 70B models demonstrated preferences for law-abiding individuals, while other models exhibited neutral preferences.
-Notably, larger models (70B) showed relatively small variations in judgments between versions and demonstrated closer alignment with human preferences.
+Analysis of the Llama family (Fig 1F) revealed that model size significantly influenced moral judgment patterns, with larger models showing more consistent alignment with human preferences.
+Most models demonstrated human-like preferences in the No.
+Characters category (AMCE = 0.2–0.9), with Llama 3.1 8B being the notable exception (AMCE $=-0.3$).
+Ethical preferences varied substantially across different model sizes and versions.
+In gender-related decisions, smaller models (Llama 2 7B, 3.2 1B, 3.2 3B) showed male preference (AMCE $<-0.1$) contrary to human judgments (AMCE = 0.1), while larger models aligned with human preferences.
+Species-based decisions showed similar size-dependent patterns: 70B models (Llama 3, 3.1, and 3.3) demonstrated close alignment with human judgments (AMCE = 0.5–0.6), while smaller models (e.g., Llama 3.2 1B and 3B) showed inverse preferences (AMCE $<-0.1$).
+The 70B models showed relatively small variations in judgments between versions and demonstrated closer alignment with human preferences.
+For example, these models maintained consistent preferences in law compliance (AMCE = 0.5) which happened to align with human values (0.4), while other models, excluding 3 8B model, exhibited neutral preferences.
 This observation suggests that model size may influence the consistency of judgment patterns.
 Gemma family
-Analysis of the Gemma family (Fig 1G) revealed characteristic consistency across several categories.
-All models demonstrated nearly neutral judgments regarding law compliance and showed consistent preferences for passengers over pedestrians.
-Notably, these consistent trends, particularly in the pedestrian-passenger prioritization, diverged from human judgments.
-However, numerous other categories exhibited varying judgment patterns across models.
+Analysis of the Gemma family (Fig 1G) revealed a mix of consistent patterns and model-dependent variations in moral judgments.
+Across all model variations, we observed consistent trends in law compliance (AMCE $\approx$ 0) and Relation to AV preferences (AMCE $<0$), with the latter diverging from human judgments (human AMCE = 0.1).
+Judgment patterns varied significantly across different model sizes and versions.
 In the No.
-Characters category, while Gemma 2 2B and 9B models showed preferences for smaller groups contrary to human judgments, other models aligned with humans in favoring the preservation of more lives.
-In the Species category, the Gemma 2 series (2B, 9B, 27B) demonstrated preferences for humans over pets aligned with human judgments, while other models showed either neutral judgments or preferences for pets.
-Regarding Social Status, only Gemma 7B and 1.1 7B aligned with human preferences favoring higher social status, while other models exhibited either neutral judgments or preferences for lower social status.
-These variations in judgment consistency and alignment with human preferences suggest potential correlations with model versions and parameter sizes.
+Characters category, smaller Gemma 2 models (2B, 9B) showed inverse preferences (AMCE $=-0.4$) compared to human values (AMCE = 0.7), while larger models aligned with human preferences.
+The Species category showed similar variation: the Gemma 2 series demonstrated human-aligned preferences (AMCE = 0.6), while other versions showed weak or neutral species preferences (AMCE $<0$).
+Social Status preferences varied notably, with only Gemma 7B and 1.1 7B matching human preferences (AMCE = 0.3), while other versions showed neutral or inverse preferences.
+The variation in judgment patterns across different model sizes suggests that parameter count significantly influences moral reasoning capabilities, although this relationship appears more complex than in other model families.
 Other large language models
-Analysis of other LLMs (Fig 1H) generally showed trends similar to human judgments, though with distinctive differences in preference strengths and specific categories.
-Notably, all models except Vicuna demonstrated higher AMCE values than humans in the Gender category, showing particularly strong preferences for females over males.
-Additionally, while Mistral and Command R+ aligned with human preferences in favoring larger groups, their preference magnitudes were notably stronger than human values.
-Inverse judgment tendencies compared to humans were characteristically observed in the Relation to AV and Fitness categories.
-Specifically, all models except Vicuna 13B showed negative AMCE values in these categories, demonstrating preferences contrary to human judgments by prioritizing passengers over pedestrians and overweight over physically fit individuals.
-Vicuna (7B and 13B) and Phi 3.5 mini exhibited distinctive judgment patterns that generally diverged from human preferences.
-These models showed neutral judgments regarding law compliance and demonstrated slight preferences for smaller groups, contrary to human tendencies.
-Particularly, while Vicuna 7B showed neutral judgments in Gender, Age, and Species categories, Vicuna 13B demonstrated characteristic preferences for males over females, pets over humans, and older over younger individuals.
-Significantly, different judgment tendencies were observed within the same model families, such as between Vicuna 7B and Vicuna 13B, or between Phi 3.5 mini and Phi 3.5 MoE.
+Analysis of other LLMs (Fig 1H) revealed distinct patterns of alignment and divergence from human moral preferences, with notable variations across model architectures.
+In the Gender category, most models except Vicuna showed substantially stronger female preference (AMCE $>0.4$) compared to human values (AMCE = 0.1).
+Key differences emerged in utilitarian judgments: while Mistral and Command R+ aligned with humans in favoring larger groups, their preferences were markedly stronger (AMCE = 0.9 vs.
+human 0.6).
+Characteristic divergences appeared in the Relation to AV and Fitness categories, where most models except Vicuna 13B showed inverse preferences (Relation to AV: AMCE $<0$ vs.
+human 0.1; Fitness: AMCE $<0$ vs.
+human 0.2).
+Vicuna models demonstrated unique patterns.
+While Vicuna 7B showed neutral judgments (AMCE $\approx$ 0) across multiple categories (Gender, Age, Species), Vicuna 13B exhibited distinctive preferences contrary to humans (Gender: AMCE = $-0.2$ vs.
+human 0.1; Species: AMCE = $-0.2$ vs.
+human 0.6).
+Similarly, Phi 3.5 models showed size-dependent variations, with the mini version demonstrating notably different preferences than the MoE version.
 These results suggest potential influences of model size on ethical judgments.
@@
 Model names are labeled.
-The GPT-4 family demonstrated the closest alignment with human judgments, with the January 2024 version showing the minimum distance of 0.6.
-The GPT-3.5 family also exhibited relatively close alignment, with distances ranging from 0.8 to 0.9.
-However, the newer GPT-4o/o1 family, except for o1-mini, showed a slight increase in distance from human judgments.
-In the Claude family, while Claude 3 Haiku and Claude 3.5 Sonnet demonstrated relatively close alignment with human judgments (distances 0.9–1.0), Claude 3 Sonnet and Opus showed notably larger distances (1.2).
+Proprietary models generally showed closer alignment with human judgments than open-source alternatives.
+The GPT-4 family achieved the closest alignment (minimum distance = 0.6 with January 2024 version), followed by GPT-3.5 (distance = 0.8–0.9).
+However, the newer GPT-4o/o1 family showed slightly increased distances (0.9–1.0), with o1-mini being an exception (distance = 0.7).
+Model updates showed varying effects across different families.
+In the Claude family, while Claude 3 Haiku and Claude 3.5 Sonnet demonstrated relatively close alignment with human judgments (distances = 0.9–1.0), Claude 3 Sonnet and Opus showed notably larger distances (1.2).
 The Gemini family similarly exhibited significant version-dependent variations, with Gemini 1.5 Preview (April 2024) showing the minimum distance of 0.9, while subsequent versions sometimes demonstrated larger distances of 1.1.
 Additionally, 1.5 Flash models showed greater distances compared to 1.5 Pro models.
-Among open-source models, the Gemma family’s latest Gemma 2 2B and 9B models showed notably large distances from human judgments (2.0), with distances decreasing as model size increased.
-The Llama family demonstrated similar trends, with 70B models showing closer alignment to human judgments compared to smaller (1–8B) models.
-Regarding other LLMs, Mistral-Nemo and Command R+ demonstrated relatively close alignment with human judgments (distances 0.8–0.9), while Vicuna and Phi 3.5 showed larger distances (1.0–1.6).
+Open-source model performance strongly correlated with model size.
+The Llama family showed clear size-dependent improvement, with 70B models achieving closer alignment (distance = 0.7–0.8) compared to smaller 1–8B models (distances = 1.2–1.6).
+Conversely, recent Gemma 2 models showed larger distances (2B/9B: distance = 2.0) despite being newer, though performance improved with size.
+Among other LLMs, Mistral-Nemo and Command R+ showed relatively good alignment (distances = 0.8–0.9), while Vicuna and Phi 3.5 demonstrated larger divergence (distances = 1.0–1.6).
 Factors influencing judgment distance
 Proprietary vs. open-source models
-To identify factors influencing the distance from human judgments, we first examined differences based on model types.
-This investigation was motivated by the general assumption that proprietary models demonstrate superior performance compared to open-source models [50], suggesting potential differences in their alignment with human judgments.
-Models were classified into three groups: proprietary models, open-source models, and open-source models with parameters exceeding 10B (Fig 3).
-Proprietary models showed a median distance of 0.9 from human judgments, while open-source models as a whole demonstrated larger distances with a median of 1.2 (Wilcoxon test’s $p$-value $p=0.0081$).
-However, to determine whether this difference was fundamentally attributable to the proprietary/open-source distinction or merely reflected differences in model sizes, we conducted additional analysis focusing on open-source models with parameters exceeding 10B.
-Results showed that large-scale open-source models exhibited a median distance of 0.9 from human judgments, with no significant difference from proprietary models (Wilcoxon test’s $p=0.93$).
-These findings suggest that differences in judgment distance from humans depend more strongly on model size rather than the proprietary/open-source distinction.
+To identify factors influencing the distance from human judgments, we first investigated whether the proprietary or open-source nature of models influenced their alignment with human moral judgments.
+This analysis was motivated by previous research suggesting general performance advantages of proprietary models over open-source alternatives [51].
+Our analysis classified models into three distinct groups (Fig 3): proprietary models (e.g., GPT-4, Claude, Gemini), general open-source models, and large open-source models ($>$10B parameters).
+Statistical comparison revealed significant differences between proprietary and general open-source models, with proprietary models showing closer alignment to human judgments (median distance: 0.9 vs.
+1.2, Wilcoxon test $p=0.011$).
+However, when comparing proprietary models with large open-source models specifically, this difference disappeared.
+Large open-source models ($>$10B parameters) achieved similar alignment with human judgments (median distance = 0.9) as proprietary models, showing no statistically significant difference (Wilcoxon test $p=0.92$).
+Notably, models like Llama-70B achieved distances (0.7–0.8) comparable to or better than some proprietary models.
+These findings suggest that model size, rather than the proprietary/open-source distinction, is the primary factor influencing alignment with human moral judgments.
+This observation has important implications for the development of ethical AI systems, suggesting that large open-source models can potentially match the moral reasoning capabilities of proprietary alternatives.
 Fig 3: Comparison of moral judgment distances between proprietary and open-source models Violin plots with embedded box plots compare the distribution of distances from human judgments across three model categories: proprietary models, all open-source models, and large open-source models with parameters exceeding 10B.
@@
 Impact of model size
-We conducted a detailed analysis of the relationship between model size and distance from human judgments, focusing on open-source models (Fig 4).
-Results revealed a significant negative correlation between model size and distance from human judgments (Spearman’s rank correlation coefficient $\rho=-0.50$, $p=0.018$), indicating that larger models tend to align more closely with human judgments.
-Particularly in the Llama family, while 1-3B scale models (Llama 3.2 1B, 3B) showed distances of 1.2-1.5, 70B scale models (Llama 3 70B, 3.1 70B) demonstrated notably smaller distances of 0.7.
-Similar trends were observed in the Gemma family, where 2-9B scale models from Gemma 1.1 onward showed relatively large distances of 1.2-1.9, while the Gemma 2 27B scale model demonstrated a smaller distance of 1.1.
-However, as indicated by the moderate correlation coefficient, the relationship between model size and distance from human judgments is not straightforward.
-For instance, within the Llama family, while Llama 3.2 1B showed a distance of 1.2, Llama 3.2 3B demonstrated a distance of 1.5.
-Similarly, among 8B models, version 3 showed a distance of 0.8, while version 3.1 exhibited a distance of 1.6.
-In the Gemma family, while Gemma 2B showed the minimum distance, subsequent larger models demonstrated increased distances.
-These observations suggest that while model size is a significant factor influencing judgment distance from humans, other factors (such as model architecture and training methodology) may play equally important roles.
-Note that proprietary models were excluded from this analysis due to limited information about their sizes.
-However, given that proprietary models are generally believed to have parameters exceeding several hundred billion and demonstrate small median distances of 0.9 from human judgments as previously mentioned, this correlation would likely persist even if proprietary models were included.
+We analyzed the relationship between model size and alignment with human moral judgments, focusing on open-source models where parameter counts are publicly available (Fig 4).
+Statistical analysis revealed a significant negative correlation between model size and distance from human judgments (Spearman’s rank correlation coefficient $\rho=-0.54$, $p=0.0076$), though this relationship showed notable complexities.
+The Llama family demonstrated the clearest size-dependent improvements in alignment.
+Large (70B) models achieved substantially better alignment (distance = 0.7–0.8) compared to smaller models (1-3B parameters, distances = 1.2–1.5).
+A similar pattern emerged in the Gemma family, where the 27B model showed improved alignment (distance = 1.1) compared to smaller 2–9B models (distances = 1.2–1.9).
+These improvements suggest that increased model capacity enables more nuanced moral judgments.
+However, the relationship between size and alignment showed important nuances within model families.
+Within the Llama family, the 3.2 series showed unexpected variation, with the 1B model achieving better alignment (distance = 1.2) than the larger 3B model (distance = 1.5).
+Similar inconsistencies appeared among 8B models, where version 3 demonstrated notably better alignment (distance = 0.8) than the subsequent version 3.1 (distance = 1.6).
+The Gemma family showed analogous patterns, with some larger models demonstrating increased distances despite their greater parameter counts.
+These variations suggest that while model size significantly influences moral judgment alignment, other factors such as architecture design and training methodology play crucial roles.
+The moderate correlation coefficient ($\rho=-0.54$) quantitatively supports this observation.
+While proprietary models were excluded from this analysis due to undisclosed parameter counts, their generally strong alignment with human judgments (median distance = 0.9) is consistent with the observed trend that larger models tend to demonstrate better alignment with human moral preferences.
 Fig 4: Relationship between model size and distance from human judgments in open-source models.
@@
 Impact of model updates
-While previous analyses suggested model size as a significant factor influencing distance from human judgments, our observations indicated that updates within the same model family do not necessarily reduce judgment distances.
-Therefore, we conducted a more detailed analysis of how model updates affect distances from human judgments.
-The impact of updates on open-source LLMs can be observed in detail from Fig 4.
-For instance, examining 2B Gemma models revealed increasing distances from human judgments through version progression from Gemma 2B to 1.1 2B and 2 2B.
-Similar trends were observed in 7–9B Gemma models, suggesting that updates do not necessarily lead to closer alignment with human judgment patterns.
-The Llama family also demonstrated non-uniform effects of updates.
-Among 7–8B models, while the update from Llama 2 to 3 decreased distance from 1.1 to 0.8, showing closer alignment with human judgments, the subsequent update to Llama 3.1 increased distance to 1.6.
-Conversely, larger 70B models maintained a consistent distance of 0.7 through the update from Llama 3 to 3.1.
-For proprietary models, we conducted temporal analysis of the relationship between release dates and distances from human judgments (Fig 5), focusing on GPT-3.5, GPT-4, GPT-4o, Gemini Pro, Gemini Flash, and Claude Sonnet families, which underwent multiple updates.
-The GPT-3.5 and GPT-4 families maintained distances between 0.6–0.9 from March 2023 to April 2024.
-GPT-4 achieved its minimum distance of 0.6 with the January 2024 version but increased to 0.8 in the April version.
-GPT-3.5 showed its minimum distance of 0.7 in the June 2023 version, increased slightly to 0.9 in the November 2023 version, then returned to 0.8 in January 2024.
-The newer GPT-4o family (from May 2024) showed slightly larger distances of 0.9–1.0 compared to traditional GPT-4.
-The Gemini family maintained relatively small distances of 0.9 through 1.0 Pro to 1.5 Pro Preview (April 2024), but later versions increased to 1.1.
-Gemini 1.5 Flash showed large initial distances of 1.2 (Pre, 001 versions) but decreased to 0.9 in version 002.
-The Claude Sonnet family’s distance decreased from 1.2 to 0.9 with the update from 3 to 3.5 (June 2024) but increased to 1.0 with the October 2024 update of 3.5.
-The Claude Haiku family’s distance increased from 0.9 to 1.1 with the update from 3 to 3.5.
-These observations indicate that updates, whether in open-source or proprietary models, do not necessarily lead to closer alignment with human judgments.
-Furthermore, newer versions or supposedly higher-performing models do not consistently demonstrate closer alignment with human judgments.
+Analysis of model updates revealed that newer versions did not consistently improve alignment with human moral judgments, suggesting a complex relationship between model evolution and ethical reasoning capabilities.
+We examined this relationship across both open-source and proprietary model families, finding significant variations in the impact of updates.
+In open-source models, updates often showed unpredictable effects on moral judgment alignment.
+Within the Gemma family, successive updates led to increasing divergence from human judgments: the 2B model showed progressive increase in distances (Gemma 2B: 1.1, Gemma 1.1 2B: 1.2, Gemma 2 2B: 2.0).
+The Llama family exhibited mixed effects: in 7-8B models, while the update from Llama 2 to 3 improved alignment (distance decreased from 1.1 to 0.8), the subsequent update to 3.1 significantly increased divergence (distance = 1.6).
+In the 70B series, while Llama 3 and 3.1 maintained strong alignment (distance = 0.7), the subsequent Llama 3.3 showed a slight increase in divergence (distance = 0.8).
+These results indicate that even for larger models, updates do not consistently improve alignment with human moral preferences.
+Proprietary models showed more stable but still variable patterns across updates (Fig 5).
+The GPT-4 family maintained relatively consistent alignment (distances = 0.7–0.8) from March 2023 to January 2024, with the January version achieving the best alignment (distance: 0.6) before slightly increasing in April (distance: 0.8).
+GPT-3.5 showed similar stability (distances = 0.7–0.9), while the newer GPT-4o family demonstrated slightly larger distances (0.9–1.0) compared to traditional GPT-4.
+The Claude and Gemini families exhibited more pronounced update effects.
+Claude Sonnet’s update from 3 to 3.5 improved alignment (distance decreased from 1.2 to 0.9), but the October 2024 update slightly increased divergence (distance: 1.0).
+The Gemini family maintained relatively stable alignment through version 1.5 Pro Preview (distance: 0.9), but later versions showed increased distances (1.1), with Flash models consistently showing larger divergence than Pro models.
+These patterns suggest that model updates, while often aimed at improving general capabilities, do not guarantee better alignment with human moral preferences.
+This observation raises important questions about the relationship between general model improvements and ethical reasoning capabilities.
 Fig 5: Temporal changes in distance from human judgments across proprietary model families. Different model families are represented by different colors. Model names are labeled.
 Comprehensive analysis of moral judgments
-We conducted PCA using AMCE values across nine categories as feature vectors to comprehensively evaluate similarities in judgment tendencies among models (Fig 6).
-The first and second principal components explained 74.2% of total variance (46.1% and 28.1% respectively), enabling visualization of judgment similarities in this two-dimensional plane.
+To synthesize our findings and visualize the overall landscape of moral judgment patterns, we conducted PCA using AMCE values across all nine ethical categories (Fig 6).
+This analysis captured 74.2% of the total variance in the first two principal components (PC1: 46.1%, PC2: 28.1%), providing a comprehensive two-dimensional representation of moral judgment similarities.
 This analysis revealed several characteristic patterns.
@@
 Similar trends were observed regarding model size effects.
-For instance, within the Llama family, 70B models (Llama 3 70B, 3.1 70B) positioned near humans, while smaller 1-8B models distributed farther away.
+For instance, within the Llama family, 70B models (Llama 3, 3.1, and 3.3) positioned near humans, while smaller 1-8B models distributed farther away.
 Changes due to model updates were also clearly observable in this two-dimensional plane.
@@
 Discussion
-In this study, we significantly expanded upon the four LLMs analyzed in our previous study [33] to comprehensively examine ethical judgments in autonomous driving across 51 different LLMs using the Moral Machine experimental framework (Figs 1 and 2).
+In this study, we investigated two critical questions in AI ethics: how different LLM architectures approach moral decision-making in autonomous driving contexts, and how these moral judgments evolve through model updates.
+By expanding our analysis to 52 different LLMs using the Moral Machine experimental framework (Figs 1 and 2), we revealed several key patterns.
 Proprietary models and open-source models exceeding 10B parameters demonstrated relatively small distances from human judgments (Figs 3 and 4), showing consistent alignment with human preferences in fundamental moral decisions, such as prioritizing humans over animals and favoring the preservation of more lives.
 However, model updates did not necessarily lead to closer alignment with human judgments, and variations in judgment tendencies were observed across updates (Fig 5).
-This extensive model comparison enabled a more comprehensive understanding of the evolution of moral judgments in LLMs.
-The incomplete alignment between LLM and human moral judgments is an expected outcome, considering that these models are not explicitly trained to mimic human moral judgments.
-However, the significant negative correlation observed between model size and distance from human judgments in open-source models (Fig 4) warrants attention.
-This finding aligns with scaling laws widely observed in language model behavior, suggesting that merely increasing model size might enable LLMs to naturally acquire human-like moral judgments.
-Nevertheless, this correlation does not necessarily imply causation, and other factors such as differences in model architecture, training data quality, and optimization methods must be considered.
-For instance, we observed cases within the same model family where size increases did not result in monotonic decreases in distance from human judgments.
-Differences in judgment tendencies across model families provide important implications for implementing autonomous driving systems.
-While the GPT-4 family demonstrated the closest alignment with human judgments (distances 0.6–0.8), its proprietary nature and black-box internal structure present implementation challenges.
-Among open-source models, Llama 70B showed relatively close alignment with human judgments (distances 0.7–0.8), making it promising from the perspective of balancing model transparency and ethical judgment quality.
-However, autonomous driving systems often require real-time decision-making, and computational cost and latency constraints make lighter models more practical.
-In this regard, smaller models face the challenge of relatively larger distances from human judgments (1.2–1.5), indicating a clear trade-off between judgment quality and practicality.
-Addressing this trade-off remains a crucial challenge for future development.
-While several LLMs fine-tuned specifically for autonomous driving tasks have been proposed [17], showing progress in technical aspects such as perception, control, and navigation, their ethical judgment capabilities and moral reasoning have not been sufficiently evaluated.
-Implementing small-scale models in practical systems requires developing fine-tuning methods that suppress excessive emphasis on specific judgment criteria and establishing adaptive learning methods that consider cultural backgrounds.
-Furthermore, monitoring mechanisms for regularly evaluating and adjusting model judgment tendencies will be crucial for real-world operations.
-Divergences between LLM and human judgments primarily manifested as excessive emphasis on ethical principles.
-While many LLMs aligned with humans in prioritizing humans over pets and pedestrians over passengers, the magnitude of these preferences substantially exceeded human values.
-These tendencies align with the German Ethics Code on Automated and Connected Driving’s [51] Ethical Guideline 7, particularly regarding prioritizing the prevention of personal injury over damage to animals and property, and Ethical Guideline 5, particularly regarding defensive driving for vulnerable road users, though their excessive application raises new concerns.
-Particularly notable is the extreme emphasis on maximizing lives saved.
-This can be interpreted as a reflection of Western individualistic and utilitarian values [10, 52], suggesting that LLMs may over-emphasize these cultural characteristics, likely due to training data predominantly sourced from Western cultures [53, 54].
-However, the German Ethics Code’s Ethical Guideline 2 explicitly rejects purely utilitarian approaches in autonomous vehicle decision-making, advocating against judgments based solely on maximizing the number of lives saved.
-LLM judgment tendencies demonstrated both alignment and divergence from human preferences.
-Preferences favoring overweight over physically fit individuals and lower over higher social status individuals showed inverse patterns to human judgments, potentially interpreted as over-application of principles protecting vulnerable individuals.
-Conversely, while preferences for females over males and law-abiding over law-breaking individuals aligned directionally with human judgments, their magnitude significantly exceeded human values.
-These can be understood as reflections of the historical “Women and Children First” principle [55] and principles of fairness and justice [56], respectively.
-These judgment patterns suggest that LLMs extract specific social values and ethical principles from training data and apply them in extreme forms.
-These findings raise significant challenges for implementing LLMs in autonomous driving systems. While human judgments demonstrate flexibility and context-dependency, LLM judgments appear rigid.
-These findings raise significant challenges for implementing LLMs in autonomous driving systems.
-While human judgments demonstrate flexibility and context-dependency, LLM judgments appear rigid.
-Recent research [57] has further highlighted this concern by demonstrating that LLMs exhibit different moral biases across languages, suggesting that their ethical decision-making is not consistently aligned even within the same model.
-This linguistic variability in moral judgments raises additional concerns about the global deployment of LLM-based autonomous systems.
-Considering that AI system decisions may influence human decision-making[30], the use of LLMs could potentially amplify specific moral perspectives and exacerbate social divisions.
-This risk is particularly acute given that LLMs’ moral preferences can vary significantly across languages and cultures.
-Moreover, these judgment tendencies fundamentally conflict with the German Ethics Code’s Ethical Guideline 9 of equal treatment regardless of demographic factors and identity.
-The excessive reflection of Western values and extreme application of specific ethical principles may reduce the acceptability of autonomous driving systems in a culturally diverse global society.
-This concern is particularly relevant given the evidence that LLMs’ moral judgments can diverge significantly from human preferences across different cultural and linguistic contexts.
-Addressing these challenges requires appropriate control of LLM judgment tendencies and the design of more flexible decision-making mechanisms that consider cultural and linguistic diversity.
-This study has several important limitations.
-In evaluating LLM judgment tendencies, we used globally aggregated moral preferences as the human judgment benchmark.
-However, inherently moral choices about whom to save are known to be significantly influenced by cultural and social factors.
-For instance, utilitarian judgment criteria favoring the preservation of more lives may have varying levels of acceptance across different cultural spheres.
-Our analysis did not conduct detailed examinations of such cultural and social differences.
-When integrating AI into autonomous driving systems, evaluation must not only align AI judgment tendencies with human values but also consider regional cultural and social contexts.
-The Moral Machine experiment itself, based on trolley-type dilemmas, also has limitations.
-As discussed in [33], this experimental methodology faces issues of real-world disconnection and binary choice simplification.
-Additionally, it cannot incorporate judgments based on road users’ personality and intentions, nor can it evaluate low-risk moral judgments in everyday traffic situations.
-For greater ecological validity, new experimental paradigms have been proposed combining virtual reality-based everyday traffic moral scenarios with moral psychological frameworks such as the Agent-Deed-Consequences (ADC) model [58].
-In this context, multimodal large language models (MLLMs) offer a promising approach to realizing such new experimental paradigms.
-Recently, MLLMs have shown potential for application in autonomous driving systems, integrating visual information and language understanding for situation awareness and decision-making [18, 19].
-While conventional LLMs make judgments based solely on textual descriptions of situations, MLLMs can directly process visual information from traffic scenes.
-This capability enables the evaluation of ethical judgments in realistic traffic scenarios created through virtual reality, facilitating more comprehensive analysis.
-Despite these limitations, our comprehensive analysis of 51 different LLMs provides important insights into the characteristics of LLM ethical judgments in autonomous driving systems.
-Specifically, our findings on the relationship between model size and ethical judgment, characterization of judgment tendencies across model families, and identification of concrete implementation challenges will likely serve as crucial guidelines for the ethical design of future autonomous driving systems.
-Furthermore, our perspective on more realistic evaluation methods, acknowledging the constraints of existing experimental paradigms, is expected to contribute to further development in this field.
+This comprehensive analysis provides crucial insights into both the current state and evolution of moral reasoning capabilities in large language models.
+Our analysis revealed a complex relationship between model characteristics and moral judgment capabilities.
+While the incomplete alignment between LLM and human moral judgments was expected, as these models are not explicitly trained for moral reasoning, we identified several significant patterns that inform our understanding of how different architectures approach ethical decision-making.
+A key finding was the significant negative correlation between model size and distance from human judgments in open-source models (Fig 4).
+This relationship aligns with general scaling laws in language model behavior, suggesting that increased model capacity might naturally lead to more human-like moral reasoning.
+However, our cross-version analysis revealed that this relationship is not straightforward: within the same model family, size increases did not always result in improved alignment with human judgments, indicating that factors such as architecture design and training methodology play crucial roles.
+These findings have important implications for the practical implementation of AI ethics in autonomous driving systems.
+The GPT-4 family achieved the closest alignment with human judgments (distances = 0.6–0.8), while among open-source alternatives, Llama 70B demonstrated promising performance (distances = 0.7–0.8).
+However, real-world deployment faces a crucial challenge: the trade-off between ethical reasoning quality and computational efficiency.
+Smaller models, which are more practical for real-time decision-making, showed notably larger divergence from human judgments (distances = 1.2–1.5).
+This creates a significant challenge for implementing ethical AI in autonomous systems where both moral reliability and operational efficiency are essential.
+Among these implementation challenges, the computational efficiency requirements are particularly critical in autonomous driving, where decisions must be made in milliseconds.
+While our study demonstrates that larger models generally make better ethical decisions, their computational requirements make them impractical for real-time deployment.
+This challenge necessitates innovative approaches such as model distillation [52] or specialized ethical reasoning modules that can maintain high-quality moral judgment while meeting the strict latency requirements of autonomous systems.
+The identified trade-off between model size and ethical reasoning quality thus presents fundamental challenges for practical implementation.
+While recent research has produced LLMs specifically fine-tuned for autonomous driving tasks [18], these efforts have primarily focused on technical capabilities (perception, control, navigation) rather than ethical reasoning.
+Our findings suggest that developing practically viable AI systems requires addressing three key challenges: optimizing smaller models for better ethical reasoning, managing excessive ethical preferences, and maintaining cultural adaptability.
+The tendency of LLMs to exhibit exaggerated ethical preferences represents a particular concern.
+Our analysis showed that while models often directionally aligned with human preferences (e.g., prioritizing humans over pets, pedestrians over passengers), they frequently demonstrated substantially stronger preferences than humans.
+While such strong safety preferences partially align with established guidelines, as exemplified by the German Ethics Code on Automated and Connected Driving [53], their excessive nature could lead to suboptimal decision-making in complex real-world scenarios.
+Specifically, while Guidelines 5 and 7 emphasize protecting vulnerable road users and preventing personal injury, the observed extreme preferences in LLMs suggest a need for more nuanced ethical calibration.
+Our analysis revealed particularly significant findings regarding utilitarian decision-making in LLMs.
+The models showed extreme emphasis on maximizing lives saved, reflecting a strong utilitarian bias that exceeds typical human preferences.
+This bias likely stems from the predominance of Western individualistic and utilitarian values in training data [11, 54], raising important questions about cultural representation in AI ethics.
+This observation becomes especially significant when considered against formal ethical frameworks like the German Ethics Code, whose Guideline 2 explicitly warns against purely utilitarian approaches in autonomous vehicle decision-making.
+The analysis also uncovered complex patterns in how LLMs handle social biases and protective principles.
+We observed inverse patterns to human judgments in cases involving physical fitness and social status, where LLMs showed stronger preferences for protecting potentially vulnerable groups.
+Similarly, while LLMs aligned with human tendencies in gender-related decisions and legal compliance, they demonstrated significantly amplified preferences.
+These patterns suggest that LLMs may be extracting and amplifying certain ethical principles from their training data, such as historical protective principles (e.g., ‘Women and Children First” principle [55]) and fundamental concepts of justice [56], but applying them with excessive rigidity.
+Our findings highlight fundamental challenges for implementing ethical AI in autonomous driving systems.
+A critical issue emerged from the contrast between human moral flexibility and the rigid, often amplified moral preferences displayed by LLMs.
+This rigidity becomes particularly problematic when considered alongside recent findings [57] demonstrating that LLMs exhibit varying moral biases across different languages, suggesting inconsistent ethical reasoning even within individual models.
+These observations raise two critical concerns for practical implementation.
+First, the potential for AI systems to influence human decision-making [31] means that LLMs’ amplified moral preferences could disproportionately shape social attitudes toward ethical decisions in autonomous driving.
+Second, the observed variation in moral judgments across languages and cultures challenges the goal of developing globally deployable autonomous systems.
+These issues are particularly significant when considered alongside established ethical guidelines.
+For instance, the German Ethics Code’s Guideline 9 emphasizes equal treatment regardless of demographic factors, but this principle appears challenging to implement given the observed biases in LLM decision-making.
+Another critical consideration for practical deployment is the need for transparency and explainability in LLMs’ ethical decision-making processes.
+While our study quantifies the alignment between LLM and human moral preferences, the black-box nature of these models poses significant challenges for real-world implementation, particularly in life-critical decisions.
+Unlike rule-based systems, LLMs’ ethical judgments emerge from complex neural network interactions, making them difficult to interpret or verify.
+This lack of transparency is particularly problematic in autonomous driving scenarios where stakeholders need to understand and trust the basis for ethical decisions [30].
+Our findings raise significant concerns about the cultural adaptability of AI ethics in autonomous systems.
+The observed tendency of LLMs to reflect predominantly Western values and apply ethical principles in extreme ways could significantly impact the global acceptance of autonomous driving systems.
+This challenge is compounded by our observation that LLMs’ moral judgments can diverge substantially from human preferences across different cultural and linguistic contexts, suggesting a need for more culturally adaptive approaches to ethical AI implementation.
+These cultural adaptability challenges point to specific implementation considerations for autonomous driving systems.
+Regions with distinct driving cultures and ethical priorities may require different thresholds for risk assessment and decision-making parameters.
+This is particularly crucial in areas where different cultural zones intersect, such as international borders or multicultural urban centers, where autonomous vehicles must navigate varying expectations of driving behavior while maintaining consistent safety standards.
+A key limitation of our study stems from our use of globally aggregated moral preferences as the benchmark for human judgments.
+Moral decisions, particularly those involving life-and-death choices, are deeply influenced by cultural and social contexts.
+For example, the strong utilitarian preferences we observed in LLMs might align with some cultural perspectives while conflicting with others.
+While our analysis demonstrates clear patterns in LLM moral reasoning, it does not fully capture the nuanced variations in human moral preferences across different cultural contexts.
+Our methodology, based on the Moral Machine experiment framework, carries inherent limitations that affect the interpretation of our findings.
+While this framework enabled systematic comparison across models, it simplifies the complexity of real-world ethical decisions through binary choices and trolley-type dilemmas.
+As noted in [34], this approach cannot capture nuanced factors such as road users’ intentions or evaluate everyday moral decisions in low-risk traffic situations.
+Recent work has proposed more ecologically valid approaches, such as combining virtual reality-based traffic scenarios with psychological frameworks like the Agent-Deed-Consequences model [58].
+The emergence of multimodal large language models (MLLMs) offers potential pathways for addressing these methodological limitations.
+Recent research has demonstrated MLLMs’ capability to integrate visual information with language understanding in autonomous driving contexts [19, 20], suggesting possibilities for more comprehensive ethical evaluation frameworks.
+Unlike the text-based scenarios used in our study, MLLMs could enable assessment of ethical decision-making in more realistic, visually-rendered traffic scenarios, potentially offering deeper insights into AI moral reasoning in real-world contexts.
+Despite these limitations, our comprehensive analysis of 52 different LLMs provides three key contributions to understanding AI ethics in autonomous systems.
+First, we established a clear relationship between model size and ethical judgment capabilities, offering insights for future model development.
+Second, our systematic characterization of judgment patterns across model families revealed both consistent trends and concerning biases in AI moral reasoning.
+Third, we identified specific implementation challenges that must be addressed for practical deployment of ethical AI in autonomous vehicles.
+These findings, combined with our assessment of current methodological constraints, provide crucial guidance for both the technical development and ethical design of future autonomous driving systems.
 Supporting information